Calculate the material turnover ratio for the yers 199 from the following details:

(mx) (my)
Opening stock 25,000 87,500
Closing stock 15,000 62,500
Purchase 1,90,000 1,25000
1
Expert's answer
2020-09-28T08:02:25-0400
"Solution"

To calculate the material turnover ratio also referred to as inventory turnover ratio, the following steps are followed.

Step 1.

Calculate material consumed for both X and Y.

"Material\ Consumed=Opening\ stock+\ Purchases\ -\ Closing\ stock.\\
\implies\ X=25000+190000-15000\\
=200,000"

Material Y will also be obtained as,

"Y=87000+125000-62500\\
=150,000"


Step 2.

Let us now ascertain the Average Inventory of each material item.

"Average\ inventory=\frac{Opening\ stock+\ Closing\ Stock}{2}\\
\implies\ X=\frac{25000+15000}{2}=20,000\\

For Y=\frac{87,500+62500}{2}=75,000"


Step 3.

Obtain the material/Inventory consumed.

"Inventory/Material\ Turnover\ Ratio\\
 = Value\ of\ Materials\ Consumed\ during\ the\ period."

This can be obtained as

"X=\frac{200,000}{20,000}=10\ Times\\
Y=\frac{150,000}{75,000}=2Times"


 In the above problem inventory/Material Turnover Ratio is highest in case of material X and lowest in case of Material "Y".